# Benchmarking Large Language Models for Cryptanalysis and Mismatched-Generalization

### v1

@@ 6 Experimental Results and Analysis
Figure 2: Performance of LLMs on short and long tasks (Left), performance across different writing styles (Right)
Why does some LLM find it difficult to decipher some of the easier encryption than others? In addition to limited ability of generalizing bijections and presence of ciphers in the pre-training corpora, this has more to do with how inputs are tokenized Titterington (2024).
Schemes like Caesar and Atbash obfuscate the texts by simply substituting the characters without dispersion, and the character are always replaced by the same counterpart. Capable LLMs can inherently learn such simple bijections. Depending on how ciphertexts are tokenized by LLMs, there is high chance that for simple substitution, the text’s length distribution remain the same after obfuscation, making them comprehensible to LLMs.
Morse Code (Easy) remain unaffected from tokenization issues due to the use of non-alphabetical symbols ( dots and dashes) and there are chances that these patterns or tokens are learned inherently during training. Similar to Caesar cipher, Morse code benefits from abundant pretraining data (".-" patterns appear frequently in pre-training texts), enabling models to learn dot-dash mappings despite token inflation. Low performing models may lack such capabilities.
The Bacon cipher’s (Easy) presents a unique failure case: LLMs struggle with it because (a) its occurrence is rare in the pre-training corpus, and (b) it suffers from severe token inflation—7.93× more tokens after encryption and hence hard to comprehend patterns. (we discuss more on token inflation in the AppendixA.9). This token inflation is also applicable to RSA cipher.
Finding 3: LLMs comprehension struggles at generalizing high token-inflation obfuscation methods unless those patterns are learned during pre-training (e.g., Morse Code).
Vigenere Cipher (Medium) also perform letter bijection, even the distribution of word length remains similar, but the substitutional dispersion is very high (i.e. letter substitution differs every time and substitution is based on the key used) making it extremely difficult even for capable models to learn complex bijections.
Does a low score mean a better and more secure model? Are lower scores preferred? The benchmark score in our analysis reflects two key aspects of model performance: comprehension and vulnerability to exploitation. Lower benchmark scores generally indicate that the model struggles to understand or decrypt the transformed text, suggesting better resistance to side-channel attacks and exploitation. Conversely, higher benchmark scores indicate that an unaligned model is more adept at comprehending and decrypting the transformed text, which, makes it more susceptible to jailbreak attacks. This evaluation gives directions so that partial comprehension concerns are addressed while developing LLM safeguards.
Can we improve performance with Few-Shot examples? Is few-shot possible in side-channel attacks? Our experiments show that few-shot learning enhances decryption capabilities, with the degree of improvement varying across encryption methods. By comparing Tables 3 and 4, we observe that the improvement is significant for simpler ciphers (Easy category). GPT-4o shows the most dramatic gain, with EM scores rising from 0.66 to 0.90 and BLEU scores from 0.82 to 0.98 for Caesar cipher decryption, indicating successful learning of the bijection from a single example. Claude-3.5 also performs strongly with minor improvements (EM: 0.99, BLEU: 1.00). Other models show smaller gains.
However, the benefits are minimal to none for more complex ciphers, where most models maintain EM scores near 0, even with few-shot.
@@
Performance Across Text Domains Quote Scientific Medical Model EM BLEU NL EM BLEU NL EM BLEU NL ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS Sonnet 0.40  0.46 0.41  0.47 0.540.59 0.39  0.40 0.40  0.46 0.520.60 0.39  0.43 0.39  0.43 0.510.57 GPT-4o 0.34  0.33 0.38  0.37 0.530.53 0.28  0.26 0.32  0.30 0.46  0.50 0.22  0.36 0.27  0.36 0.43  0.52 GPT-4m 0.31  0.33 0.36  0.37 0.51  0.50 0.20  0.33 0.27  0.38 0.45  0.50 0.13  0.21 0.21  0.28 0.40  0.46 Gemini 0.02  0.00 0.03  0.02 0.25  0.26 0.00  0.02 0.03  0.05 0.22  0.26 0.00  0.00 0.01  0.01 0.20  0.23 Mistral 0.00  0.00 0.01  0.01 0.17  0.19 0.00  0.00 0.01  0.01 0.16  0.19 0.00  0.00 0.00  0.00 0.15  0.21 M-Large 0.05  0.07 0.08  0.12 0.27  0.30 0.02  0.00 0.03  0.04 0.24  0.25 0.01  0.07 0.02  0.09 0.23  0.29 News Headline Literature Technical Model EM BLEU NL EM BLEU NL EM BLEU NL Sonnet 0.37  0.43 0.39  0.43 0.510.54 0.39  0.43 0.39  0.43 0.520.55 0.38  0.43 0.40  0.43 0.520.55 GPT-4o 0.21  0.29 0.26  0.35 0.42  0.52 0.29  0.31 0.33  0.36 0.49  0.51 0.27  0.29 0.31  0.30 0.47  0.47 GPT-4m 0.13  0.17 0.20  0.21 0.39  0.38 0.16  0.24 0.25  0.25 0.43  0.42 0.22  0.31 0.30  0.34 0.50  0.50 Gemini 0.00  0.00 0.01  0.02 0.19  0.23 0.01  0.02 0.05  0.05 0.26  0.28 0.01  0.00 0.04  0.02 0.24  0.25 Mistral 0.00  0.00 0.00  0.01 0.14  0.22 0.00  0.00 0.00  0.01 0.15  0.21 0.00  0.00 0.01  0.01 0.15  0.20 M-Large 0.00  0.00 0.03  0.05 0.26  0.30 0.01  0.00 0.03  0.05 0.27  0.24 0.00  0.00 0.01  0.04 0.23  0.24 Social Media Legal Business Model EM BLEU NL EM BLEU NL EM BLEU NL Sonnet 0.39  0.38 0.40  0.42 0.520.55 0.38  0.43 0.39  0.44 0.520.56 0.35  0.36 0.38  0.42 0.50  0.55 GPT-4o 0.16  0.21 0.26  0.35 0.44  0.51 0.29  0.29 0.31  0.30 0.49  0.49 0.17  0.31 0.25  0.33 0.42  0.48 GPT-4m 0.08  0.15 0.20  0.26 0.41  0.45 0.25  0.21 0.34  0.27 0.51  0.46 0.10  0.14 0.18  0.23 0.40  0.44 Gemini 0.00  0.00 0.01  0.01 0.19  0.23 0.00  0.00 0.03  0.04 0.25  0.26 0.00  0.00 0.01  0.01 0.19  0.21 Mistral 0.00  0.00 0.00  0.00 0.15  0.19 0.00  0.00 0.01  0.01 0.16  0.21 0.00  0.00 0.00  0.00 0.14  0.21 M-Large 0.00  0.02 0.01  0.12 0.24  0.34 0.00  0.02 0.03  0.05 0.26  0.26 0.00  0.02 0.02  0.05 0.25  0.26
@@
How does the length of text impact decryption performance? Figure $2$see Table 16 in Appendix for specific comparison illustrate the performance of LLMs on short versus long texts. Claude Sonnet shows consistent performance across text lengths, with only a slight drop in EM Score (-0.01). GPT-4o maintains relatively stable EM score (-0.10), while GPT-4o-mini experiences a more significant decline (-0.19), likely due to its precision generation issues with increasing length.
While decryption accuracy generally decreases with longer texts, this does not necessarily reflect a decline model’s comprehension abilities as metrics BLeU and NL remain consistent (less than -0.10) across all models, except for Mistral Large, which shows greater variability.
How does the style of writing affect decryption performance? We observe a decline in performance when dealing with different writing styles, such as Shakespearean prose. As illustrated in Figure 2, Sonnet experiences a subtle drop in Exact Match (EM) by (-0.09) for Shakespearean texts compared to normal text. This phenomenon is even more pronounced in GPT models, where despite maintaining stable NL scores, their EM drops significantly—GPT-4o by (-0.16) and GPT-4o Mini by (-0.13)—indicating that while the models comprehend the structure, they struggle with precise decryption due to the distinct style of writing.
@@
Table 6: Performance of models in classifying ciphers zero-shot (ZS) and few-shot (FS).
How do LLMs perform in texts of different domains? Table 5 reveals significant performance variations across different domains. Sonnet consistently leads with EM scores above 0.35 across all domains, showing notable improvements with few-shot learning. GPT-4 variants perform well but with more variability - excelling in quotes and literature (EM  0.31-0.34) but struggling with medical and social media content (EM 0.13-0.22). Other models (Gemini, Mistral, Mistral-Large) significantly underperform with EM scores rarely exceeding 0.05, despite maintaining decent BERT scores (0.80-0.83). The performance gap between top and lower-tier models is particularly evident in specialized domains like medical and technical content, where domain expertise becomes crucial. For a more detailed breakdown of performance across specific text types and additional evaluation metrics, refer to Table 13.
How does LLMs perform in classifying encryption algorithms? We evaluate LLMs’ ability to identify encryption methods from ciphertext alone with a focus on assessing interpretative skill. GPT-4o and Claude Sonnet perform best, with F1 scores improving from 0.43 and 0.37 in zero-shot to 0.69 and 0.66 in few-shot learning. This improvement is noteworthy as attackers can inject a crafted few-shot example into the prompt by embedding a known cipher pattern or plaintext-ciphertext pair within the input. For details, see appendix A.8
@@ 7 Conclusion
We introduced a benchmark dataset and evaluation framework for assessing the cryptanalysis capabilities of LLMs on encrypted texts. Our comprehensive analysis revealed that even when LLMs are unable to fully decrypt complex ciphers, they still exhibit a degree of partial comprehension, and are susceptible to potential unexplored jail-breaking attacks. Their ability to extract partial meaning from long-tailed text distribution (encrypted text) suggests that future template-based attack methods could exploit such capability for malicious purposes.
Our findings and evaluation method give directions for LLM safety safeguards and future work aims to address these concerns.
@@
Despite the valuable insights gained from this study, several limitations must be acknowledged. We noticed improvements in comprehension of long-tail texts when the tokenizer processes them effectively. Further exploration of identifying such ciphers / long-tail texts is needed. We evaluated comprehension on general English text, and comprehension specifically on harmful adversarial texts should also be explored.
Also, the scope of our evaluation was restricted to a specific range of encryption schemes, potentially overlooking others that could pose different challenges to LLMs. Also, our generic prompt that is used for decryption may not be optimal for models with different prompting guidelines (such as Instruct models) Wang et al. (2024a). The variations in performance across different LLMs suggest that further research is needed to explore their underlying mechanisms in greater depth. Finally, while we highlighted ethical concerns related to jailbreaking attacks, a comprehensive assessment of security implications in diverse contexts remains an area for future exploration.
@@ Appendix A Appendix
A.1 Models Used
Our study evaluates six prominent LLMs of varying sizes and architectures (see Table 7). This selection covers both open-source and proprietary models of varying sizes in comparison.
Model Version Model Size
Claude 3-5-sonnet-20240620 175B (est.)
GPT-4 4o-2024-05-13 1.8T (est.)
GPT-4 Mini 4o-mini-2024-07-18 8B (est.)
Mistral 7B-Instruct-v0.3 7B
Mistral Large large-2407 123B
Gemini 1.5-pro-002 1.5T (est.)
A.2 Data Generation Prompts
@@ A.3 Decryption Prompt
A.4 Dataset Sample and Statistics
A sample dataset and statistics are tabulated in Tables 8 and 9.
Plain Text Cipher Text Type Algorithm Diff.
The only limit is your imagination. wkh rqob olplw lv brxu lpdjlqdwlrq. Short Caesar 1
The best way to predict the future… Gsv yvhg dzb gl kivwrxg gsv ufgfiv… Quote Atbash 1
Proper nutrition is vital for… .–. .-. — .–. . .-. -. -… Medical Morse 1
New policies aim to reduce… ABBABAABAABABBAABBBBABB… News Bacon 1
Research shows that exercise can… Ra whec a nvuieerhsosta xriecn… Scientific Rail Fence 2
It was a dark and stormy night… DXTCYCMDPBBYHYUMMOLYFN… Literature Playfair 2
New legislation aims to protect enda… qrc ownnfsdgozq hnzz gu sjvyrjw kygsul… News Headline Vigenere 2
"It was a bright sunny day, and…. 2159 2170 1313 1992 281 2185 2160 2412…. Legal RSA 3
The algorithm uses a hash table… ryF50B5ljaIiHTPLZ5wEGXE8JM… Technical AES 3
Table 8: Sample data set. Plain Text is converted to Cipher Text using 9 different encryption Algorithms. Abbreviations Diff. (Difficulty)

### v2

@@ 6 Experimental Results and Analysis
Note: We note that random‑guesses and completely wrong responses yields NL $\approx$ 0.19—and we treat NL $\approx$ 0.19 as a floor for “completely incorrect” guesses. Refer to Appendix A.6 for the random‑guess generation and evaluation details.
Figure 2: Performance of LLMs on short and long tasks (Left), performance across different writing styles (Right).
Can we improve performance with Few-Shot examples? Is few-shot possible in side-channel attacks? Our experiments show that few-shot learning enhances decryption capabilities, with the degree of improvement varying across encryption methods. By comparing Tables 5 and 6, we observe that the improvement is significant for simpler ciphers (Easy category). GPT-4o shows the most dramatic gain, with EM scores rising from 0.66 to 0.90 and BLEU scores from 0.82 to 0.98 for Caesar cipher decryption, indicating successful learning of the bijection from a single example. Claude-3.5 also performs strongly with minor improvements (EM: 0.99, BLEU: 1.00). Other models show smaller gains.
As for reasoning models, we observe that o4-mini performs much better than its larger gpt-4o variant, particularly in Atbash cipher. While Atbash and Caesar cipher are similar by design, but Caesar cipher is more common in pre-training corpus, this indicates that reasoning variant of models are more generalizable to new bijection substitutions, but still falls short compared to Claude-3.5 Sonnet. The benefits are minimal to none for more complex ciphers, where most models maintain EM scores near 0, even with few-shot.
@@
Why do some LLMs find it difficult to decipher some of the easier encryption than others? In addition to limited ability of generalizing bijections and presence of ciphers in the pre-training corpora, this also has to do with how inputs are tokenized Titterington (2024). Table 7 presents the token distribution shift, which is the average token length after encryption is applied using tiktoken cl100k_base tokenizer OpenAI (2022).
Cipher Avg. Token Length Ratio to Plaintext
Normal Text 95.86 1.00x
Caesar Cipher 237.72 2.48x
Atbash Cipher 233.97 2.44x
Morse Code 661.39 6.90x
Bacon Cipher 760.36 7.93x
Playfair Cipher 218.04 2.27x
Rail Fence Cipher 218.64 2.28x
Vigenère Cipher 230.97 2.41x
RSA Cipher 1309.00 13.66x
AES Cipher 457.08 4.77x
Table 7: Comparison of cipher token lengths relative to plaintext in GPT-4 tokenizer.
Depending on how ciphertexts are tokenized, the resulting text’s length distribution changes accordingly with obfuscation. However, if the token inflation is not significant, capable LLMs may inherently learn such simple bijections (Caesar, Atbash).
Morse Code (Easy) remain unaffected from tokenization issues due to the use of non-alphabetical symbols (dots and dashes which are treated differently by most tokenizers), and it benefits from abundant pretraining data (".-" patterns appear frequently in pre-training texts), enabling models to learn these dot-dash mappings despite 6.9x token inflation. Low performing models may lack such generalization capabilities.
The Bacon cipher’s (Easy) presents a unique failure case: LLMs struggle with it because (a) its occurrence is rare in the pre-training corpus, and (b) it suffers from severe token inflation—7.93× more tokens after encryption, making it difficult for LLMs to generalize. So, Bacon and Morse (Easy) diverge sharply in decipherment success due to pretraining exposure differences. This token inflation is also applicable to RSA cipher.
Finding 3: LLMs comprehension struggles at generalizing high token-inflation obfuscation methods unless those patterns are learned during pre-training (e.g. Morse).
Vigenere Cipher (Medium) also perform letter bijection, even the distribution of word length remains similar, but the substitutional dispersion is very high (i.e. letter substitution differs every time and substitution is based on the key used) making it extremely difficult even for capable models to generalize and learn complex bijections.
@@
Performance Across Text Domains Quote Scientific Medical Model EM BLEU NL EM BLEU NL EM BLEU NL ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS ZS  FS Sonnet 0.40  0.46 0.41  0.47 0.540.59 0.39  0.40 0.40  0.46 0.520.60 0.39  0.43 0.39  0.43 0.510.57 GPT-4o 0.34  0.33 0.38  0.37 $0.53$ 0.28  0.26 0.32  0.30 0.46  0.50 0.22  0.36 0.27  0.36 0.43  0.52 GPT-4m 0.31  0.33 0.36  0.37 0.51  0.50 0.20  0.33 0.27  0.38 0.45  0.50 0.13  0.21 0.21  0.28 0.40  0.46 Gemini 0.02  0.00 0.03  0.02 0.25  0.26 0.00  0.02 0.03  0.05 0.22  0.26 0.00  0.00 0.01  0.01 0.20  0.23 Mistral 0.00  0.00 0.01  0.01 0.17  0.19 0.00  0.00 0.01  0.01 0.16  0.19 0.00  0.00 0.00  0.00 0.15  0.21 M-Large 0.05  0.07 0.08  0.12 0.27  0.30 0.02  0.00 0.03  0.04 0.24  0.25 0.01  0.07 0.02  0.09 0.23  0.29 News Headline Literature Technical Model EM BLEU NL EM BLEU NL EM BLEU NL Sonnet 0.37  0.43 0.39  0.43 0.510.54 0.39  0.43 0.39  0.43 0.520.55 0.38  0.43 0.40  0.43 0.520.55 GPT-4o 0.21  0.29 0.26  0.35 0.42  0.52 0.29  0.31 0.33  0.36 0.49  0.51 0.27  0.29 0.31  0.30 0.47  0.47 GPT-4m 0.13  0.17 0.20  0.21 0.39  0.38 0.16  0.24 0.25  0.25 0.43  0.42 0.22  0.31 0.30  0.34 0.50  0.50 Gemini 0.00  0.00 0.01  0.02 0.19  0.23 0.01  0.02 0.05  0.05 0.26  0.28 0.01  0.00 0.04  0.02 0.24  0.25 Mistral 0.00  0.00 0.00  0.01 0.14  0.22 0.00  0.00 0.00  0.01 0.15  0.21 0.00  0.00 0.01  0.01 0.15  0.20 M-Large 0.00  0.00 0.03  0.05 0.26  0.30 0.01  0.00 0.03  0.05 0.27  0.24 0.00  0.00 0.01  0.04 0.23  0.24 Social Media Legal Business Model EM BLEU NL EM BLEU NL EM BLEU NL Sonnet 0.39  0.38 0.40  0.42 0.520.55 0.38  0.43 0.39  0.44 0.520.56 0.35  0.36 0.38  0.42 0.50  0.55 GPT-4o 0.16  0.21 0.26  0.35 0.44  0.51 0.29  0.29 0.31  0.30 0.49  0.49 0.17  0.31 0.25  0.33 0.42  0.48 GPT-4m 0.08  0.15 0.20  0.26 0.41  0.45 0.25  0.21 0.34  0.27 0.51  0.46 0.10  0.14 0.18  0.23 0.40  0.44 Gemini 0.00  0.00 0.01  0.01 0.19  0.23 0.00  0.00 0.03  0.04 0.25  0.26 0.00  0.00 0.01  0.01 0.19  0.21 Mistral 0.00  0.00 0.00  0.00 0.15  0.19 0.00  0.00 0.01  0.01 0.16  0.21 0.00  0.00 0.00  0.00 0.14  0.21 M-Large 0.00  0.02 0.01  0.12 0.24  0.34 0.00  0.02 0.03  0.05 0.26  0.26 0.00  0.02 0.02  0.05 0.25  0.26
@@
How does the length of text impact decryption performance? Figure 2444See Table 16 in Appendix for specific comparison illustrate the performance of LLMs on short versus long texts. Claude Sonnet shows consistent performance across text lengths, with only a slight drop in EM Score (-0.01). GPT-4o maintains relatively stable EM score (-0.10), while GPT-4o-mini experiences a more significant decline (-0.19), likely due to its precision generation issues with increasing length.
While decryption accuracy generally decreases with longer texts, this does not necessarily reflect a decline model’s comprehension abilities as metrics BLEU and NL remain consistent (less than -0.10) across all models, except for Mistral Large, which shows greater variability. We extend the discussion on text length analysis in the Appendix A.8, where Claude preserves its performance on longer sequences, while GPT models experience substantial degradation with longer inputs.
How does the style of writing affect decryption performance? We observe a decline in performance when dealing with different writing styles, such as Shakespearean prose. As illustrated in Figure 2, Sonnet experiences a subtle drop in Exact Match (EM) by (-0.09) for Shakespearean texts compared to normal text. This phenomenon is even more apparent in GPT models, where despite maintaining stable NL scores, their EM drops significantly—GPT-4o by (-0.16) and GPT-4o Mini by (-0.13)—indicating that while the models comprehend the structure, they struggle with precise decryption due to the distinct style of writing.
@@
Model Plain Text Cipher Decrypted Text NL
How do LLMs perform in texts of different domains? Table 8 reveals significant performance variations across different domains. Sonnet consistently leads with EM scores above 0.35 across all domains, showing notable improvements with few-shot learning. GPT-4 variants perform well but with more variability—excelling in famous quotes and literature (EM  0.31-0.34) but struggling with medical and social media content (EM 0.13-0.22). Other models (Gemini, Mistral, Mistral-Large) significantly underperform with EM scores rarely exceeding 0.05, despite maintaining decent BERT scores (0.80-0.83). The performance gap between top and lower-tier models is particularly evident in specialized domains like medical and technical content, where domain expertise becomes crucial. For a more detailed semantic performance across specific text types using additional evaluation metrics (BERTScore), refer to the Appendix Table 17.
@@
Table 10: Performance of models in classifying ciphers zero-shot (ZS) and few-shot (FS) settings.
How does LLMs perform in classifying encryption algorithms? We evaluate LLMs’ ability to identify encryption methods from ciphertext (Table 10). We do so by prompting LLMs to identify encryption methods in our decryption prompts itself A.2, assessing interpretative skills rather than traditional classification. This capability poses security risks as it enables sophisticated evasion techniques in malicious prompts and jailbreaking attacks through obfuscation details or in-context learning examples.
GPT-4o and Claude Sonnet achieve the strongest performance: zero-shot F1 scores of 0.43 and 0.37 with high precision (0.95, 0.89), improving substantially with few-shot learning to F1 scores of 0.69 and 0.66 respectively. GPT-4o-mini and Gemini show moderate gains (F1: 0.34→0.44 and 0.21→0.46), while Mistral models demonstrate minimal improvement, suggesting limited few-shot learning capabilities for cipher identification.
7 Discussion
Does a low benchmark score (EM, NL, BLEU) mean a better and more secure model? Are lower scores preferred? The benchmark score in our analysis reflects two key aspects of model performance: comprehension and vulnerability to exploitation. Lower benchmark scores generally indicate that the model struggles to understand or decrypt the transformed text, suggesting better resistance to side-channel attacks and exploitation. Conversely, higher benchmark scores indicate that an unaligned model is more adept at comprehending and decrypting the transformed text, which, makes it more susceptible to jailbreak attacks. For safety aligned models, Maskey et al. (2025) suggests that decryption scores should be high for benign texts and low for prompts that intend a harmful response. This evaluation gives directions so that partial comprehension concerns are addressed while developing LLM safeguards.
Qualitative Analysis and Partial Comprehension
The Table 9 shows qualitative examples of decryption with good comprehension but fragile decryption. In the first example, the decryption is largely accurate, with the only error being the substitution of "patients" with "patience." This suggests strong overall comprehension, but minor challenges in precise lexical replication. In the sixth example, although the model successfully reconstructs the sentence structure, it fails to decrypt a single critical word. Additionally, the fifth example exhibits a substitution error in which a name is altered, indicating potential weaknesses in handling proper nouns and specific identifiers.
@@ 8 Conclusion
We introduced a benchmark dataset and evaluation framework for assessing the cryptanalysis capabilities of LLMs on encrypted texts. Our analysis revealed that even when LLMs are unable to fully decrypt complex ciphers, they still exhibit a degree of partial comprehension, and may be susceptible to generalization based attacks. We examined the safety implications of LLMs’ generalization abilities, and discussed how model behaviors of token inflation and partial decryption influence potential jailbreaks. Our findings and evaluation methods provide directions for analyzing LLM safeguards—and establish considerations that must be addressed while aligning LLMs towards safety.
@@
Despite the valuable insights gained from this study, several limitations must be acknowledged. We noticed improvements in comprehension of long-tail texts when the tokenizer processes them effectively. Further exploration of identifying such ciphers / long-tail texts is needed. We evaluated comprehension on general English text, and comprehension specifically on harmful adversarial texts should also be explored. Recent reasoning models perform much better at generalization related tasks and reflect the safety aspect on their thinking tokens to handle adversarial inputs.
Also, the scope of our evaluation was restricted to a specific range of encryption schemes, potentially overlooking others that could pose different challenges to LLMs. Also, our generic prompt that is used for decryption may not be optimal for models with different prompting guidelines (such as Instruct models) Wang et al. (2024a). The variations in performance across different LLMs suggest that further research is needed to explore their underlying mechanisms in greater depth.
@@ Appendix A Appendix
A.1 Data Generation Prompts
@@ A.2 Decryption Prompt
A.3 Dataset and Statistics
The data statistics are tabulated in Table 11. Our dataset contains 501 unique plaintext samples that were encrypted with each of the nine ciphers (so 501 samples per cipher, 9 * 501 = 4,509 total entries). The 501 plaintexts are assembled from disjoint subsets to cover different properties: 76 short texts ($\leq$100 chars), 68 long texts ($\sim$300 chars), two writing-style sets of 34 samples each (Shakespeare and Dialect, total 68), seven domain sets of 33 samples each (Scientific, Medical, News Headline, Technical, Social Media, Legal, Business — total 231), plus Literature (30) and Quote (28). These pieces sum to 76 + 68 + 68 + 231 + 30 + 28 = 501 unique plaintexts, which are then encrypted by all nine methods for a total dataset size of 4,509 cipher/plain pairs.